OpenStudy (anonymous):

Can someone explain the steps to get the answer to this: 3q= 8-7p 8p-18=q

OpenStudy (campbell_st):

well this can be solved by substitution, take the 2nd equation and substitute it into the 1st equation 3(8p - 18) = 8 - 7p or 24p - 54 = 8 - 7p now add 7p to both sides of the equation 31p -54 = 8 I'll let you finish the solution of p. when you have p, substitute it into either of the initial equations to find q... hope it helps

OpenStudy (jiteshmeghwal9):

\[7p+3q=8\]\[8p-q=18\]First of all make the coefficient of any of the same variables equal\[=>7p+3q=8\]\[3(8p-q)=(18)3\]\[=>7p+3q=8\]\[24p-3q=54\]now eliminate '3q'. therefore,\[31p=62\]now find 'p' & then put the value of 'p'in any of the equation toget 'q':)

OpenStudy (jiteshmeghwal9):

my method is the elimination method & @campbell_st has used the substitution method. so i think u will nt gt confused @AmberLove01 :)

OpenStudy (anonymous):

I have been stuck on this question for a while && I have finally understood it! I found the elimination one a little bit easier, but thanks for both helping me.

OpenStudy (jiteshmeghwal9):

yw :)